# More for Keys, Less for Values: Adaptive KV Cache Quantization

### v1

@@ Abstract
This paper introduces an information-aware quantization framework that adaptively compresses the key-value (KV) cache in large language models (LLMs). Although prior work has underscored the distinct roles of key and value cache during inference, our systematic analysis—examining singular value distributions, spectral norms, and Frobenius norms—reveals, for the first time, that key matrices consistently exhibit higher norm values and are more sensitive to quantization than value matrices. Furthermore, our theoretical analysis shows that matrices with higher spectral norms amplify quantization errors more significantly. Motivated by these insights, we propose a mixed-precision quantization strategy, KV-AdaQuant, which allocates more bit-width for keys and fewer for values since key matrices have higher norm values. With the same total KV bit budget, this approach effectively mitigates error propagation across transformer layers while achieving significant memory savings. Our extensive experiments on multiple LLMs (1B - 70B) demonstrate that our mixed-precision quantization scheme maintains high model accuracy even under aggressive compression. For instance, using 4-bit for Key and 2-bit for Value achieves an accuracy of 75.2%, whereas reversing the assignment—2-bit for Key and 4-bit for Value—yields only 54.7% accuracy. The code is available at https://tinyurl.com/kv-adaquant.
More for Keys, Less for Values: Adaptive KV Cache Quantization
Mohsen Hariri1, Lam Nguyen1, Sixu Chen1, Shaochen Zhong2, Qifan Wang3, Xia Hu2, Xiaotian Han1, Vipin Chaudhary1 1Case Western Reserve University; 2Rice University; 3Meta mohsen.hariri@case.edu
@@ 1 Introduction
Figure 1: Key cache needs more bits. (Left): Spectral norms of the key cache (blue) and value cache (orange) across layers in Llama3.3-70B show that key caches consistently exhibit higher norms. (Right): GSM8k accuracy for two schemes: $\text{K}_{2}\text{V}_{4}$, representing 2-bit allocation for the K cache and 4-bit allocation for the V cache and $\text{K}_{4}\text{V}_{2}$, representing 4-bit allocation for the K cache and 2-bit allocation for the V cache, demonstrates that allocating more bits to the key cache maintains strong performance, confirming the efficacy of norm-aware, mixed-precision quantization.
Large language models (LLMs) have advanced NLP by enabling parallel processing and capturing complex dependencies Vaswani et al. (2017); Sutskever et al. (2014), which have scaled from GPT-1’s 117M Radford et al. (2018), GPT-2’s 1.5B Radford et al. (2019) to GPT-3’s 175B Brown et al. (2020) and GPT-4’s 1.8T parameters OpenAI et al. (2024). Open-source models like Llama have grown to 405B Touvron et al. (2023), with Mistral Large (123B) Mistral-AI (2023) and DeepSeek V3 (671B) DeepSeek-AI (2023) also emerging. In parallel with increasing parameter counts, longer context windows are becoming essential for chain-of-thought reasoning, necessitating an expansion of the KV cache and contributing to memory bottlenecks Wei et al. (2022). New systems support context windows of up to 1M tokens DeepMind (2025), 200k tokens OpenAI (2024, 2025), and 130k tokens DeepSeek-AI et al. (2025). NVIDIA’s Blackwell architecture introduces low-precision formats like FP8, FP4, and INT4 NVIDIA (2025), emphasizing the need for precise quantization methods. The KV cache is a critical component in these models and its effective operation is paramount and significant.
To accommodate larger context windows, the KV cache can become a significant memory bottleneck. One effective solution is to quantize the key and value caches from high-precision formats (e.g., FP32 or BF16) to lower-precision ones (e.g., INT4 or INT2). However, we first discovered that the Key cache generally exhibits higher norms than the Value cache, making it more sensitive to uniform quantization. Since matrices with larger norms are more prone to quantization errors, bit-width allocation should be tailored to each matrix’s norm distribution. To confirm this, we conducted experiments with Llama3.3-70B (see Figures 1 (Left) and 2), as the Key cache (blue) consistently shows elevated norm values compared to the Value cache (orange). Furthermore, Figure 1 (Right) demonstrates that the K(4bit)V(2bit) scheme significantly outperforms the K(2bit)V(4bit) configuration on GSM8K dataset. These results underscore the need for a norm-aware, mixed-precision strategy that allocates higher precision to Keys and lower precision to Values.
Guided by these inspiring observations, we propose a mixed-precision quantization strategy, KV-AdaQuant, which allocates different bit-width to more bits for keys and fewer bits for values. since key matrices consistently have higher norm values. With the same total KV bit budget, this approach effectively mitigates error propagation across transformer layers while achieving significant memory savings. KV-AdaQuant uses singular values, spectral norms, and Frobenius norms to assign more bits to Keys and fewer to Values, balancing memory savings with preserving critical information. This work makes the following contributions: Guided by these insights, we propose a mixed-precision quantization strategy, KV-AdaQuant, which assigns a higher bit-width to key caches and a lower bit-width to value caches—reflecting the consistently higher norm values observed in key matrices. KV-AdaQuant leverages metrics such as singular values, spectral norms, and Frobenius norms to identify the critical information in KV cache. By maintaining the same overall KV bit budget, our approach effectively mitigates error propagation across transformer layers while offering significant memory savings. Our contributions include:
  • We systematically investigate the KV cache in transformers, uncovering structural patterns in how information is stored and retrieved.
  • We introduce KV-AdaQuant, an information-aware quantization method that leverages singular value distributions to allocate mixed precision—assigning higher bit-widths to keys and lower bit-widths to values.
  • We demonstrate the efficacy of our approach on the GSM8K benchmark, achieving substantial memory savings without compromising performance.
2 Key and Value are Different
This section investigates the differences between key and value caches to enhance quantization strategies.
2.1 Singular Value Distribution of the KV Cache
An extensive collection of models, ranging in size up to 70B parameters (Llama 3.3 70B Grattafiori et al. (2024)), is examined to explore the singular values of the KV cache. Random batches are sampled from the C4 Dodge et al. (2021), GSM8K Cobbe et al. (2021), and MMLU Hendrycks et al. (2021) datasets, and KV caches are extracted across all attention heads and layers. Figure 3 illustrates the distribution of singular values for the KV cache, starting from the 5th largest singular value onward, highlighting how keys and values diverge. Additionally, Figure 7 in Appendix C presents the full-range distribution, further revealing the discrepancy in larger singular values. Beyond individual singular values, the spectral norm ($\|M\|_{2}$) and the Frobenius norm ($\|M\|_{F}$) are computed, providing empirical evidence of the substantial gap in key and value matrix singular values and illustrating how this characteristic of the KV cache can inform the theoretical analysis of quantization error propagation in multi-layer attention networks. This discrepancy underscores the importance of careful precision allocation to mitigate performance degradation while saving more memory in the inference time.
2.2 Rank of KV Cache
The rank of a matrix corresponds to the number of nonzero singular values and determines the effective dimension of its column or row space. In multi-head self-attention, the key ($K$) and value ($V$) matrices possess one fixed dimension, $d_{\mathrm{head}}$, and one variable dimension corresponding to the sequence length, $\mathrm{seq\_len}$. Consequently, $K$ and $V$ can be analyzed in two distinct regimes: (a) when $\mathrm{seq\_len}<d_{\mathrm{head}}$, and (b) when $\mathrm{seq\_len}\geq d_{\mathrm{head}}$. In the first regime, the matrices are relatively short and thus offer limited opportunities for compression. Accordingly, many existing KV-cache compression methods either avoid compressing in this stage or employ specialized strategies that differ from those used later.
When $\mathrm{seq\_len}\geq d_{\mathrm{head}}$, the $K$ and $V$ matrices assume rectangular shapes with dimensions $\mathrm{seq\_len}\times d_{\mathrm{head}}$. Although the exact rank is data-dependent (since $K=XW_{K}$, where $X$ is drawn from a high-dimensional embedding space), empirical observations suggest that both key and value cache typically approach a rank close to $d_{\mathrm{head}}$. Therefore, in practical settings it is reasonable to assume that there are approximately $d_{\mathrm{head}}$ nonzero singular values for both $K$ and $V$. In the remainder of this section, we focus on this regime to analyze how spectral and Frobenius norm considerations guide our quantization strategy.
2.3 Stability of the Dequantized KV Cache: Error Propagation
The spectral norm of a matrix $M$, denoted $\|M\|_{2}$, quantifies the maximum stretching factor that $M$ can impart on any vector. In a multi-layer transformer (e.g., Llama 3.3 Grattafiori et al. (2024) with 80 layers or Phi-4 Abdin et al. (2024) with 16 layers), the sequential application of transformations implies that quantization errors introduced at one layer can propagate and amplify in subsequent layers.
To illustrate this effect, consider a simplified layer without normalization or additional scaling Elhage et al. (2021), where the hidden state at layer $(l+1)$ is given by
$h_{l+1}=h_{l}+W_{l}h_{l},$ (1)
Let $W_{l}$ denote the weight matrix at layer $l$. Define $\widetilde{W}_{l}$ as the result of quantizing the KV cache, and $W_{l}$ as the unquantized version. Then, the local error incurred at layer $l$ can be bounded as
$\begin{split}\|h_{l+1}-\widetilde{h}_{l+1}\|_{2}&=\|(W_{l}-\widetilde{W}_{l})h_{l}\|_{2}\\ &\leq\|W_{l}-\widetilde{W}_{l}\|_{2}\|h_{l}\|_{2}.\end{split}$ (2)
Propagating such errors through multiple layers yields a cumulative effect that is proportional to the product of the spectral norms at each layer. In other words, even small perturbations introduced by quantization can be significantly amplified if the weight matrices exhibit large singular values. Empirical observations indicate that key matrices consistently have higher spectral norms than value matrices. This suggests that errors in key matrices are more prone to propagate and magnify throughout the transformer stack. As a result, it is advisable to employ less aggressive quantization (e.g., 4-bit rather than 2-bit) for key matrices, while value matrices—characterized by lower spectral norms—can tolerate more aggressive quantization without incurring substantial distortion.
2.4 Total Energy of KV Cache (MSE Perspective)
The Frobenius norm of a matrix, denoted by $\|M\|_{F}$, is defined as the square root of the sum of the squared entries:
$\|M\|_{F}=\sqrt{\sum_{i,j}M_{i,j}^{2}}$
Equivalently, it can be viewed as the square root of the sum of the eigenvalues of $M^{\top}M$, meaning it captures the total “energy” of $M$. In the context of mean-squared error (MSE), the Frobenius norm provides a direct indicator of how much overall distortion might occur when quantizing the matrix.
Quantization Error and the Frobenius Norm.
When a matrix $M$ is quantized to a lower-precision representation $\widetilde{M}$, the quantization error in MSE terms is often measured by the squared Frobenius norm:
$\|M-\widetilde{M}\|_{F}^{2}=\sum_{i,j}(M_{i,j}-\widetilde{M}_{i,j})^{2}.$
A common assumption is that this error scales with both the range of the entries in $M$ and the granularity determined by the bit precision. Formally, one may write
$\mathbb{E}\bigl{[}\|M-\widetilde{M}\|_{F}^{2}\bigr{]}\;\propto\;\|M\|_{F}^{2}\times 2^{-2b},$
where $b$ denotes the number of bits used per entry. [^1] The proportionality depends on the quantizer design and the distribution of entries in $M$. For more details see Appendix A. For matrices with the same dimensions (e.g., key and value cache), the MSE and squared Frobenius norm are directly proportional by a factor of $1/mn$. Consequently, when analyzing quantization error for matrices of identical size:
  • Minimizing $\|M-\widetilde{M}\|_{F}^{2}$ is equivalent to minimizing $\text{MSE}(M,\widetilde{M})$, as the proportionality constant $1/mn$ does not affect optimization or relative comparisons.
  • The scaling relationship $\mathbb{E}\bigl{[}\|M-\widetilde{M}\|_{F}^{2}\bigr{]}\propto\|M\|_{F}^{2}\times 2^{-2b}$ holds equivalently for MSE, since both metrics share the same dependence on $b$ and $\|M\|_{F}^{2}$.
Thus, the Frobenius norm of the quantization and mean square of the quantization error can be used interchangeably for key and value cache, as they encode the same error structure up to a constant scaling factor.

### v2

# Quantize What Counts: Bit Allocation Insights Informed by Spectral Gaps in Keys and Values

@@ Abstract
Large Language Models (LLMs) have introduced significant advancements to the capabilities of Natural Language Processing (NLP) in recent years. However, as these models continue to scale in size, memory constraints pose substantial challenge. Key and Value cache (KV cache) quantization has been well-documented as a promising solution to this limitation. In this work, we provide two novel theorems aimed at enhancing KV quantization methods. Our first theorem, termed Key-Value Norm Disparity, states that the key weight matrices by nature carry richer information compared to the value weight matrices, as evidenced by higher spectral and Frobenius norms across most of the layers. Our second theorem, Key-Driven Quantization, posits that prioritizing the quantization precision of keys over values induces significant improvements to the overall quantization performance. In particular, assigning greater precision to the keys compared to the values achieves a higher degree of precision reduction with minimal impact on model accuracy. We validate these theorems through theory and extensive experiments on several state-of-the-art LLM architectures and benchmarks. These findings offer valuable guidelines for improving KV cache quantization strategies, facilitating more efficient memory utilization without compromising model performance across diverse NLP tasks. Source code is available at https://github.com/mohsenhariri/spectral-kv.
@@ 1 Introduction
LLMs have introduced significant advancements to the capabilities of NLP via their ability to express and capture intricate ideas [1, 2], and a key reason for these advancements is model scaling. GPT has seen a tremendous increase across its four versions: GPT-1 at 117M [3], GPT-2 at 1.5B [4], GPT-3 at 175B [5], and GPT-4 at 1.8T parameters [6]. In the open-source space, Llama has grown to 2T [7], Mistral Large is now 123B [8], and DeepSeek V3 is at 671B [9].
However, with this growth in model size comes a prohibitive memory bottleneck in the KV cache [10]. An increase in parameter count requires an increase in context-window length to accommodate chain-of-thought reasoning, which in turn expands the KV cache central to the attention unit. That expansion can impose a restrictive memory bottleneck on the LLM. Already, supported context windows in SOTA systems have grown significantly [11, 12, 13, 14], underscoring the need for a solution. One powerful remedy is quantization, wherein high-precision numbers are reduced to low-precision formats [15, 16]. In LLMs, this can be applied to the KV-cache tensors (e.g., from BF16 to INT4) yielding substantial memory savings with minimal performance loss.
@@
Figure 1: Keys carry greater importance. (Left): Keys exhibit higher norms than values across all Llama 3.3-70B layers. (Right): GSM8K accuracy is higher under $\mathrm{K}_{2}\mathrm{V}_{4}$ (2-bit keys, 4-bit values) than $\mathrm{K}_{4}\mathrm{V}_{2}$.
This process is known as KV cache quantization. Extensive work has been done in designing novel KV quantization frameworks, yet none so far have provided deep investigation into mixed-precision schemes, in particular those which focus on the bit allocations split between the Key and Value caches. Instead, previous works have treated the bit allocation as a hyperparameter, seeing it as an object for optimization via grid search rather than taking a deeper dive into the properties of Keys and Values. We provide further motivation for the study in the form of a full description of the current literature, which is detailed in Section 2.
Therefore, there is a present need for a deep exploration into the inherent differences between the Key and the Value caches, as well as a change to the way that quantization is performed. This study aims to fill this need by uncovering fundamental theoretical ideas to improve upon KV cache quantization methods. A preview of these results is illustrated in Figure 1 (left), showing that the spectral norms of the Keys consistently exceed those of the values. Additionally, Figure 1 (right) demonstrates that accuracy across four different models is notably higher for the K4V2 quantization configuration compared to the K2V4 configuration. Our main contributions are summarized as follows:
  • We prove that the expected values of the spectral norm and Frobenius norm are consistently higher for Key weight matrices compared to Value weight matrices (Key-Value Norm Disparity theorem), implying that Key matrices inherently carry more information. We provide detailed empirical validation through observational studies on multiple LLMs.
  • Building on the Key-Value Norm Disparity, we introduce the Key-Driven Quantization theorem, positing that assigning higher bit precision to Keys over Values significantly enhances quantization efficiency. Specifically, this approach allows for greater precision reduction while maintaining minimal accuracy loss.
  • We substantiate the Key-Driven Quantization theorem through two primary forms of evidence: (1) a theoretical analysis explicitly linking it to the previously established Key-Value Norm Disparity, and (2) comprehensive empirical evaluations involving multiple SOTA LLMs, three datasets, and two quantization frameworks.
Section 2 addresses gaps in the existing literature on KV cache quantization, followed by Section 3.1, which introduces the Key-Value Norm Disparity theorem and its supporting evidence. Section 3.2 presents the Key-Driven Quantization theorem along with its corresponding proof, and Section 4 reports the empirical evaluation of both theorems. Finally, Section 5 summarizes the main contributions and discusses their limitations.
2 Related Work
KV Quantization Background.
Quantization frameworks can be categorized in several ways. One way is the time at which the quantization is performed, which is grouped into training-time and post-training (PTQ) quantization techniques [17]. Training-time quantization frameworks can struggle with the usual costs of neural network training, including the dependence on labelled data and longer training times, but these frameworks can achieve more accurate performance due to targeting the weights or activations during the training process. In contrast, post-training quantization techniques are the reverse, struggling with sometimes less accurate performance due to dealing with the values after the training has been completed but achieving lower computation overhead and no longer relying on labelled data which may not be available [16].
There are several settings for PTQ: weight-only, weight-activation, and KV quantization. Weight-only quantization focuses on the weights and achieves strong performance but can struggle with memory overhead [18, 19, 20]. Weight-activation quantization focuses on both the weights and the activations and avoids memory overhead but can struggle with achieving high accuracy [21, 22, 23]. In contrast, KV quantization offers the memory savings seen in weight-activation methods while also maintaining the performance of weight-only quantization [24]. This is due to the KV cache memory overhead inflation caused by longer sequences [25].
KV quantization frameworks can be further categorized into three groups based on the exact quantization scheme being used for the keys and values: outlier redistribution, fixed-precision, and mixed-precision [26]. In outlier redistribution, the outliers in the cache are moved elsewhere or smoothed out. Several outlier redistribution frameworks have been built to varying levels of success [22, 20, 23]. In fixed-precision, the same number of bits is used to quantize certain parts of the cache while in mixed-precision, different numbers of bits are used. Fixed-precision frameworks have also been built to varying levels of success, though they all suffer from a similar problem of neglecting the inherent properties within the KV cache as well as the varying importance of tokens [27, 28, 29].
More focus has been devoted to the third type of KV quantization framework, mixed-precision, but as yet, these works have not looked much into asymmetric allocation into keys and values [30, 24, 31, 32, 33, 34, 35]. KVQuant, for example, focuses on defining an outlier for each vector, such as channel or token, to achieve what the authors term dense-and-sparse quantization. Only KVTuner, QAQ, and SKVQ discuss the asymmetric allocation in some detail, which is described below.
KV Quantization Gaps.
As described above, extensive work has been done in designing new KV quantization frameworks, yet none so far have conducted deeper investigation into mixed-precision schemes. In particular, no studies have focused on making a bit allocation split between the key matrix and the value matrix, nor have they provided a deep theoretical explanation for the inherent differences between the two matrices. Only three studies to date have touched upon this area, with none providing explanations for potential differences between the key and value matrix importance.
The closest study has been a KV quantization framework called KVTuner, which employs layerwise quantization based on the sensitivity of the layers to the KV cache quantization mode and precision pairs [35]. The study makes a brief mention of the difference in attention score errors and perplexity score errors, observing that precision pairs with more bits allocated for keys than for values achieves similar accuracy than for even-split pairs, such as K8V4 versus K8V8. However, the study makes no further analysis into this observation, providing no reason for why this difference is observed. In addition, the quantization approach does not use any underlying insight on this difference between key and values, instead opting for a straightforward grid search over a limited amount of precision pairs such as K8V4, K4V8, K4V2, and so on. Therefore, this allocation split of bits is not the core focus of the framework, with the focus instead being on deploying a layerwise quantization scheme.
Two other studies also touch upon the key-value difference, but like KVTuner, do not provide much depth into the difference. One of these studies is a KV quantization framework called SKVQ, which deploys a sliding window filter along the input tokens to remove tokens that fall outside the window frame and that meet certain criteria [33]. The quantization strategy achieves 2 bit key quantization and 1.5 bit value quantization, but the authors do not explore the possible reasons behind why this difference in bit allocation yielded such powerful results, instead arriving at this bit allocation via hyperparameter search.
The other study is a KV quantization framework called QAQ, which uses the attention output to perform the key quantization and the self-attention output to perform the value quantization [32]. Unlike SKVQ and all other KV quantization strategies, QAQ is the first to target the keys and values in different approaches. However, the difference in the approach lies in the type of data being used to perform the quantization, not in the exact number of bits being allocated to one versus the other. Finally, both SKVQ and QAQ are post-training quantization approaches, making them susceptible to overhead issues during the attention computation.
Figure 2: Frobenius norm for Llama Family. The x-axis represents the layer index in the model while the y-axis represents the Frobenius norm magnitude. The norms are higher for the Key weights than for the Value weights across layers.
3 Dynamics of the Norms of Key and Value
3.1 Key-Value Norm Disparity
Before formalizing our finding, a brief examination is presented on how the magnitudes of the key and value projection matrices influence KV cache quantization. In the self-attention mechanism, the key weights $W^{k}$ scale the token representations that enter the cache, whereas the value weights $W^{v}$ scale the representations that are read out once the attention scores are formed. Because quantization noise is approximately proportional to the dynamic range of the signal, the relative sizes of $\|W^{k}\|$ and $\|W^{v}\|$ determine which part of the cache is more sensitive to a given bit budget. Computing norms across four Llama-3 models (Fig. 2) show that the Frobenius norm of $W^{k}$ consistently exceeds that of $W^{v}$ in almost every layer; the same ordering holds for the spectral norm. Intuitively, this systematic gap motivates the formal statement captured in Theorem Key-Value Norm Disparity below.
Theorem 1 (Key-Value Norm Disparity).
The expectation of the Frobenius norm for the key weight matrices are greater than that for the value weight matrices.
$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}\bigr{]}\;>\;\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}\bigr{]}.$ (1)
The proof of Theorem 1 is deferred to Appendix B, but its core argument can be sketched as follows: Starting from a standard Xavier initialization[36], the analysis tracks how stochastic-gradient descent updates the Frobenius norms of $W^{k}$ and $W^{v}$. Under isotropic-input assumptions, the cross-term $\langle W^{m},\,\partial\mathcal{L}/\partial W^{m}\rangle_{F}$ becomes negligible, so the expected norm growth is governed by the squared gradient magnitudes. A careful decomposition of those gradients shows that the Key weight update contains an additional factor of the query weight matrix $Q$, whose own norm increases during training. Consequently, the expected squared gradient of $W^{k}$ dominates that of $W^{v}$, causing $\mathbb{E}\bigl{[}\|W^{k}\|_{F}\bigr{]}$ (and, by monotonicity, the spectral norm) to outpace its value counterpart. This asymptotic divergence establishes the persistent norm gap asserted Key-Value Norm Disparity.
Although Fig. 2 illustrates this disparity for the Llama family, supplementary plots for Mistral models in Appendix D confirm that the same pattern holds across diverse architectures [8]. The next subsection investigates how this norm disparity influences quantization error, laying the groundwork for data-driven bit-allocation strategies.
3.2 Key-Driven Quantization
Theorem Key-Value Norm Disparity establishes a persistent norm gap between the Key and Value weight matrices. Since quantization noise scales with the dynamic range of the signal, weight matrices of larger norm inevitably introduce larger absolute errors under identical bit precision. Consequently, assigning equal quantization budgets to the Key and Value caches is sub-optimal.
Consider an additive-residual Transformer block (Layer Norm omitted for clarity [37]),
$h_{l+1}=h_{l}+W_{l}h_{l},$
and let $\Delta_{l}$ denote the quantization perturbation applied to $W_{l}$. A single block incurs
$\|h_{l+1}-\tilde{h}_{l+1}\|_{2}=\|\Delta_{l}h_{l}\|_{2}\leq\|\Delta_{l}\|_{2}\,\|h_{l}\|_{2}.$
After $L$ layers the worst-case deviation satisfies
$\|h_{L}-\tilde{h}_{L}\|_{2}\;\leq\;\Bigl{(}\prod_{i=1}^{L}\|W_{i}\|_{2}\Bigr{)}\;\max_{1\leq i\leq L}\|\Delta_{i}\|_{2}.$
From  Key-Value Norm Disparity:
$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}^{2}\bigr{]}.$ (i)
Hence any distortion injected on the key path is amplified more strongly throughout the layers.
Let $X\in\mathbb{R}^{\mathrm{seq}\times d_{\mathrm{model}}}$ be the hidden-state matrix at a given layer. The key and value caches are produced by the same input but different projections:
$K=X\,W^{K},\qquad V=X\,W^{V}.$
Because $X$ is common to both paths, multiplying (i) on the left by $X$ and using the sub-multiplicativity of the Frobenius norm, $\|AB\|_{F}\leq\|A\|_{F}\,\|B\|_{2}$, yields
$\mathbb{E}\!\bigl{[}\|K\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|V\|_{F}^{2}\bigr{]},$ (ii)
i.e. the norm gap persists in the caches themselves.
Quantization error vs. Frobenius norm.
For a matrix $M\in\mathbb{R}^{m\times n}$ the squared Frobenius norm $\|M\|_{F}^{2}$ equals the total energy $\sum_{i,j}M_{ij}^{2}$, or equivalently the sum of eigenvalues of $M^{\top}M$ [^1] Because $\operatorname{tr}(AB)=\operatorname{tr}(BA)$, one has $\operatorname{tr}\!\bigl{(}M^{\top}M\bigr{)}=\operatorname{tr}\!\bigl{(}MM^{\top}\bigr{)}$. The non-zero eigenvalues of $M^{\top}M$ and $MM^{\top}$ are identical and equal to the squared singular values $\sigma_{i}^{2}$ of $M$; hence $\|M\|_{F}^{2}=\sum_{i}\sigma_{i}^{2}=\operatorname{tr}\!\bigl{(}M^{\top}M\bigr{)}$.. When $M$ is quantized to $\tilde{M}$ with $b$ bits, uniform-scalar quantizers yield:
[…]
where [^2] Big-$\Theta$ notation: $f(n)=\Theta(g(n))$ iff $\exists\,c_{1},c_{2}>0,\;n_{0}$ such that $c_{1}g(n)\leq f(n)\leq c_{2}g(n)$ for all $n\geq n_{0}$. the hidden constants depend only on the quantizer variant. Since Key and Value caches share the same shape, minimizing mean square error reduces to minimizing $\|\,\cdot\,\|_{F}^{2}$. When $\|K\|_{F}\gg\|V\|_{F}$ and equal bit-widths are used,
$\mathbb{E}\!\bigl{[}\|K-\tilde{K}\|_{F}^{2}\bigr{]}\;\gg\;\mathbb{E}\!\bigl{[}\|V-\tilde{V}\|_{F}^{2}\bigr{]},$
making the key cache the dominant source of degradation, formally:
Theorem 2 (Key-Driven Quantization).
Let $(b_{K},b_{V})$ denote the bit allocations assigned to the key and value caches, respectively, under a symmetric uniform-scalar quantizer. For any pair $(b_{K},b_{V})$ and $b_{K}>b_{V}$, the expected inference accuracy is strictly higher than for the swapped allocation $(b_{V},b_{K})$, provided that $\mathbb{E}\!\bigl{[}\|K\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|V\|_{F}^{2}\bigr{]}$.